connect omeworh Chapter 13 1428 poiets s Dance Studios is considering the purchase of new soud quipment that wl hance the popuanity of its aerk answers using the toe and ncial ciator menods dancing The eqpmert wlcost 526,300. Dabby is not sure how many members he new squipmen 9,800 a. What is the expected value of the cash flow? The value you compute will apply to each of the five years b. what is the expected net present value? Negative amount should be indicated by a-n sign·Do not round intermediate calculations and round your answer to 2 decimal places.) e Shouid Debby buy the new equipmeni
Answer #1

(a) Expected Value of Cash Flows = Σ CFiPi, where CFi is the cash flow with probability Pi

=> Expected Value of Cash Flows = 3890*0.2 + 5190*0.3 + 7550*0.4 + 9800*0.1 = $6335

(b)

CF0 = -26300
CF1 = 6335
CF2 = 6335
CF3 = 6335
CF4 = 6335
CF5 = 6335

NPV = -CF0 + CF1/(1+r) + CF2/(1+r)2 + ....

=> NPV = -26300 + 6335/1.12 + 6335/1.122 + 6335/1.123 + 6335/1.124 + 6335/1.125

=> NPV =$ -3463.74

(c) Since NPV is negative, there is no benefit in buying the new equipment. Hence. Debby should not buy the new equipment.